OpenStudy (anonymous):

Prove that sec2x=(1/2)(secx)(cscx)

OpenStudy (anonymous):

How far have you gotten?

OpenStudy (anonymous):

According to my calculations, that statement is not true.

OpenStudy (anonymous):

are you sure?? LS= 1/cos2x RS= 1/2sinxcosx

OpenStudy (anonymous):

Wait. Is the right side (1/2)(secx)(cscx) or 1/2sinxcosx

OpenStudy (anonymous):

okokok its sec2x=1/2(secx)(cscx)

OpenStudy (anonymous):

so RS i did 1/2(1/cosx)(1/sinx)

OpenStudy (anonymous):

In the end, that will simplify to sin(2x) = cos(2x), which can only equal at a specific value of x.

jimthompson5910 (jim_thompson5910):

If you graph the two sides separately as different functions, you'll notice they do not line up perfectly (see attached image). So that visually confirms this equation is not an identity.

OpenStudy (paxpolaris):

maybe it shoould read:\[ \color{red}\csc2x=\frac12\sec x \cdot \csc x\]

OpenStudy (anonymous):

ok thank you all!

jimthompson5910 (jim_thompson5910):

If PaxPolaris is correct with that equation, then it is an identity because graphing the two has them lining up perfectly. One graph is perfectly on top of the other as you can see in the attached gif file.

jimthompson5910 (jim_thompson5910):

Of course, this isn't proof your teacher is looking for, but it's handy to have visual confirmation.